OpenStudy (anonymous):

i need to find the perimeter of the shaded region, but first, how can I find the diameter/radius? is AE=EO?

OpenStudy (anonymous):

is the diameter=20

ganeshie8 (ganeshie8):

can u draw it out so we may calculate diameter quick ? :)

OpenStudy (perl):

@ganeshie8 can you help me with the question i posted earlier

OpenStudy (perl):

why is i^i^i^i^i^i^e= 1 ?

ganeshie8 (ganeshie8):

ugh im no better than u... i gave up that problem yesterday itself :'(

OpenStudy (anonymous):

wait up..

OpenStudy (p0sitr0n):

10pi+40 = perimeter

OpenStudy (mathmale):

S-I-M: Really, I don't want to call you "stupidinmath," so would you provide a positive nickname that I could use in stead? Referring to the drawing you've already shared: construct a right triangle EOF. The base of this triangle, EO, is half of DC, or 5. Its height is 10. What's the hypotenuse of this triangle? Twice that is your diameter.

OpenStudy (anonymous):

 the file i posted is bettter

OpenStudy (anonymous):

P0sitr0n really? c:

OpenStudy (anonymous):

mathmale, haha, nah i'm really a stupid in math .. but yeah, thanks :)

OpenStudy (mathmale):

No soap. Reallyl do want another nickname (a positive one) by which I can call you. When you start thinking of yourself as "good in math," you're a lot more likely to be good in math. So, pretty please: a new nickname. Do my comments on your math problem make sense? If so, please calculate the hypotenuse, OD (or DO).

OpenStudy (anonymous):

Oh, alright c: mathgenius then? haha. wait let me try

OpenStudy (mathmale):

I wouldn't say AE=EO without proof.

OpenStudy (mathmale):

So you go right ahead, mathgenius, and see whether you can find the hypotenuse.

OpenStudy (anonymous):

125's the hypotenuse

OpenStudy (anonymous):

so.. diameter is 20, right?

OpenStudy (mathmale):

I see you're using the Pyth. Them. Good. 5^2 + 10^2 = hyp^2 =125. Now what is the hyp?

OpenStudy (mathmale):

\[5^{2}+10^{2}=125=h ^{2}\] That's great. You've had it right; that was the radius of your semicircle. But then you deleted what you'd typed. Repeat it, please.

OpenStudy (mathmale):

If h^2=125, h=+(what?)

OpenStudy (anonymous):

√125.. 5√5

OpenStudy (mathmale):

Again, perfect, mathgenius! I'd bet that you can now solve the rest of the problem yourself. But I'm sticking around in case you have further questions.

OpenStudy (anonymous):

So.. need to multiply it by 2 to get the diameter. so.. diameter's 10√5 right?wrong?

OpenStudy (mathmale):

Let's concentrate on being right and pay little or no attention to the possibility of being wrong. To learn, we have to make mistakes. Yes, you have the diameter correct in 10Sqrt(5). Let's read the problem statement again, carefully. What's your understanding of the goal we're aiming to reach?

OpenStudy (mathmale):

Or, better phrased, what's our goal in this problem?

OpenStudy (anonymous):

Alright, thanks. finding the peri of shaded region

OpenStudy (anonymous):

how will i get that..

OpenStudy (mathmale):

Right. I'm assuming that that means the DARK region. Have you worked with "arc length" before? As in \[s=r*(\theta)\] ??

OpenStudy (anonymous):

some parts of it

OpenStudy (anonymous):

I guess that's kinda easy that part

OpenStudy (mathmale):

OK, mathgenius. We can make this problem a lot less daunting by breaking it up into smaller parts. I ask y ou to focus first on the pie-shaped figure AOD. The curved boundary is called an "arc." We have to determine the length of this arc. Sound familiar?

OpenStudy (anonymous):

Yeap, got that:)

OpenStudy (anonymous):

using s=r∗(θ) right?

OpenStudy (mathmale):

If you've had trigonmetry, it should be relatively easy for you to determine the angle EOD. Think you can do that? Once you have this central angle in degrees, you must convert it to radians. Or, better yet, obtain the central angle directly in rads. Yes, use s=r(theta). You're a mathgenius.

OpenStudy (anonymous):

EOD= 45 degrees?

OpenStudy (mathmale):

I'd prefer not to say Yes or No, but rather to ask you to demonstrate how you obtained that.

OpenStudy (anonymous):

Oh alright.. E = 90 D= 90/2 = 45 180 - (90 + 45) = 45 = EOD

OpenStudy (mathmale):

Have you had trigonometry? If so, does the following sound familiar?\[\tan \theta=\frac{ opp }{ adj }\]

OpenStudy (anonymous):

yep, oh right.. wait

OpenStudy (mathmale):

Are you familiar with the inverse trig functions, e. g., \[\theta =\tan ^{-1} \frac{ a }{ b}\]

OpenStudy (anonymous):

yes, let me find it :))

OpenStudy (anonymous):

is 63.43 right? or calculated it wrong?

OpenStudy (mathmale):

mathgenius, if you'll look carefully at triangle EOD, you'll see that the side opposite the angle EOC is 10 and the side adjacent to EOD is 5. You could do either of the following: Write "angle EOD"=\[\tan ^{-1}\frac{ 10 }{ 5}=\tan ^{-1}2,\]

OpenStudy (mathmale):

or calculate the actual value (as you have done). Your result is in degrees, so before I respond I'm going to ask you to convert that angle measurement to radians.

OpenStudy (anonymous):

1.106?

OpenStudy (mathmale):

On my calculator, \[\tan ^{-1}2=1.107 radians, or 63.435 \]degrees. So, you were RIGHT the first time.

OpenStudy (mathmale):

You're doing JUST FINE!!

OpenStudy (anonymous):

Haha finally:)

OpenStudy (mathmale):

Now please find arc length AD. I'd like to see your work, not just the answer.

OpenStudy (anonymous):

To find arc length.. just double the angle measurement right? If <AED =90 so.. Arc AD=180 am i right?

OpenStudy (mathmale):

Hold a moment. I'd like to draw a diagram for you.

OpenStudy (mathmale):